The Ghana Statistical Services in 2019 fiscal year estimated that Ghanaians saved 20 percent of their disposable income; however, due to overconsumption in the previous year, autonomous saving for 2019 was negative 100 million Ghana cedis. The government of Ghana imposed a lump sum tax of Gh ₡ 20 million. In addition to the lump sum tax the government also took 20 percent of all incomes earned in the economy in the form of proportional tax. The total government expenditure was GH ₡50 million in 2019. Total investment recorded for 2019 was 66 million.

Determine the tax function for this economy.


1
Expert's answer
2021-09-03T08:09:21-0400

Economy is in equilibrium where Aggregate Expenditure = Aggregate Income .

"AE = C + I + (G-T)"

therefore ,

"Y = C + I + (G-T)"

Given :

S = 20% Yd  , where Yd = disposable income 

Yd = Y - tY  ( where t is proportional tax )

Sa= - 100 million , (Autonomous Saving ) in 2019 

T = 20 million , (Lump sum taxes)

t = 20%Y  (Proportional taxes )

I = 66 million 

G = 50 million (Government Expenditure )


a)

Taxes = T + t  ( Autonomous saving + Proportional taxes )

Tax function = 20 million + 0.2Y